(Assignment 25 (Chapter 19) Nuclear Fission and Fusion < 3 of 10 > Review | Constants Periodic Table - Part A Certain heavy n<Assignment 25 (Chapter 19) Nuclear Fission and Fusion < 3 of 10 > Review | Constants Periodic Table Certain heavy nuclei und<Assignment 25 (Chapter 19) Problem 19.78 4 of 10 > Review Constants Periodic Table Part A How much energy is released (in kJPlease show all work

thank you!
Answer #1

Fustan fission Beryllium-7 Helium-3 Nitrogen-15 Uranium-231 Curium. 243 Thorium - 227 530 39 e 235 után → 129 g + 10ty + 21hEnergy released = 0.00233x931 Mev - 2.16923 Mer IMeV = 16x10-13 = 2.16923%/-6X16135 = 3.47*15-16 kJ